OpenStudy (anonymous):

Write an equation for the line tangent to the circle x2 + y2 = 17, at the point (4, 1).

OpenStudy (anonymous):

eq. of tangent at(x1,y1) toa circle \(x^2+y^2=a^2\) is \( xx1+yy1=a^2\)

OpenStudy (anonymous):

It has to the at the point of (4,1)

OpenStudy (anonymous):

x1=4,y1=1

OpenStudy (anonymous):

\(a^2=17\)

OpenStudy (anonymous):

it has to be written in form: y=mx+b

OpenStudy (anonymous):

so it is 4x +1y=17 second method \(find dy/dx at (4,1)\) then eq. of tangent is y-y1=m(x-x1)

OpenStudy (anonymous):

okay i got the answer..Its y=-4x+17

OpenStudy (anonymous):

can you help me on one more question..